OpenStudy (anonymous):

The lengths of plate glass parts are measured to the nearest 1/10 mm. The lengths are discretely and uniformly distributed with values at every 1/10th mm, starting at 590.0 thru 590.9 mm. Determine the mean & variance of lengths.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@ybarrap @yama_aryayee

OpenStudy (ybarrap):

It says uniformly distributed are you sure its geometric?

OpenStudy (anonymous):

I really do not know but I guess it should be done with poisson method

OpenStudy (ybarrap):

For uniform, the mean is $$ \bar{x}=\cfrac{590.0+590.9}{2} $$ The variance is $$ \sigma=\sqrt{\cfrac{n^2-1}{12}} $$ Where n=10.

OpenStudy (anonymous):

n is 9 or 10 since there are 9 portions

OpenStudy (anonymous):

between 590 and 590.9

OpenStudy (ybarrap):

There are 10 portions: 590.0 590.1 590.2 590.3 590.4 590.5 590.6 590.7 590.8 590.9

OpenStudy (anonymous):

and mean is the mean value which mean a number in the middle of these two, so do not u think it should be 590.45?

OpenStudy (anonymous):

yes

OpenStudy (ybarrap):

590.45 is correct

OpenStudy (anonymous):



OpenStudy (anonymous):

and I think variance and mean should be equal ,,

OpenStudy (ybarrap):

 BTW, variance is $$ \sigma^2=\cfrac{n^2-1}{12} $$ NOT $$ \sigma=\sqrt{\cfrac{n^2-1}{12}} $$